How many quarts of a 20% salt solution must be mixed with 15 quarts of a 35% salt solution to obtain a mixture that is 25% salt?
Answer by CharlesG2(834)

How many quarts of a 20% salt solution must be mixed with 15 quarts of a 35% salt solution to obtain a mixture that is 25% salt?
% * amount mixture A + % * amount mixture B = % * amount mixture C
20% * x + 35% * 15 = 25% * (x + 15)
0.20x + 0.35 * 15 = 0.25x + 0.25 * 15
0.20x + 5.25 = 0.25x + 3.75 (rearrange)
1.5 = 0.05x
30 quarts = x